Erin’s age is 3 times Warren’s. in 4 years she will be twice as old as he will be. How old is each?
Answer by Earlsdon(6287)

Start by assigning variables the the unknown quantities.
Let E = Erin's present age.
Let W = Warren's present age.
Erin's age (E)is (=) three times Warren's (3W), so you can write:
1) E = 3W
In four years, Erin's age (E+4) will be (=) twice Warren's (2(W+4)), so you can write:
2) E+4 = 2(W+4) Substitute from equation 1) E = 3W and solve for W.
3W+4 = 2(W+4) Simplify.
3W+4 = 2W+8 Subtract 2W from both sides.
W+4 = 8 Subtract 4 from both sides.
W = 4 This is Warren's present age.
E = 3W
E = 3(4)
E = 12 This is Erin's present age.
Check:
E+4 = 2(W+4)
12+4 = 2(4+4)
16 = 2(8)
16 = 16